OpenStudy (anonymous):

A stone is thrown upward from ground level. The initial speed is 48 feet per second. How high will it go? Without using integrals lol.

OpenStudy (anonymous):

s = ut + 0.5at^2 48(1) + 0.5(-10)(1) 48 + 5 = 43 assuming g = -10

OpenStudy (anonymous):

why is g= -10?

OpenStudy (anonymous):

@miteshchvm need to use -32.2ft/s/s because units are in feet.

OpenStudy (anonymous):

ya -32.2 you need to use g = -32.2 as it is going against gravity

OpenStudy (anonymous):

Kinematics . . . \[y_{final}=y_{initial}+v_{initial}(t)-16.1(t)^2\]

OpenStudy (anonymous):

i cant tell which is which. :(

OpenStudy (anonymous):

*from the problem

OpenStudy (anonymous):

If you want some practice with your calculus, take that equation and differentiate once with respect to 't' to get the velocity, and differentiate again to get acceleration.

OpenStudy (anonymous):

"A stone is thrown upward" = initial velocity is positive. ("The initial speed is 48 feet per second.") "from ground level." = initial height = 0. "How high will it go?" = solve for y_final.

OpenStudy (anonymous):

how do i find t? or is it the same as speed..?

OpenStudy (anonymous):

There are probably some other simplifying assumptions you will need. One: neglecting air resistance, the parabolic trajectories of objects in free-fall are symmetric, which means they take the same amount of time to go up as they do to come down.

OpenStudy (anonymous):

@cliffsedge what is y final? \[y_{final}=y_{initial}+v_{initial}(t)-16.1(t)^2 \]

OpenStudy (anonymous):

You can also differentiate the position equation to get the velocity equation and realize that at the top of the flight, velocity =0.

OpenStudy (anonymous):

@CliffSedge i guess you have written a wrong equation.

OpenStudy (anonymous):

That is never the wrong equation. It is the fundamental kinematics equation, all others are derived from it. Try this: take that equation, set it equal to zero for y_final and solve for t (use quadratic formula).

OpenStudy (anonymous):

Granted, there are short-cut equations you can use for this sort of thing, but understanding where they come from will help out in the long run.

OpenStudy (anonymous):

When you get more experience with this sort of thing, you'll see that if you differentiate the position equation: \[\frac{d}{dx}(\Delta y = vt+0.5at^2) \rightarrow \Delta v = at\] It becomes the velocity equation; if you solve it for t and substitute that in to the position equation, you'll get \[\Delta (v^2)=2a \Delta y,\] which lets you solve things without needing to know the time variable.